### v2

# Probing dark matter interactions with a RES-NOVA prototype cryogenic detector

Abstract
We report on the operation of a 13 g PbWO4 crystal, grown from archaeological Pb and operated as a cryogenic calorimeter in an underground environment. Read out with a Ge thermistor, the detector achieves a low energy threshold and, for the first time, enables the derivation of a dark matter exclusion limit using PbWO4 as target material, for both spin–dependent interactions on neutrons and spin–independent interactions. Although limited in mass and not representative of the final RES–NOVA detector design, this prototype demonstrates effective control of mechanical vibrations and low–energy noise in a cryogenic system, which is a key requirement for rare–event searches. The experiment therefore provides a proof of principle for the RES–NOVA detection concept, validating the use of archaeological Pb–based PbWO4 crystals, low–background operation, and robust data–analysis procedures. These results establish a solid technological and methodological foundation for future RES–NOVA detectors employing larger target masses and advanced thermal readout technologies.
1 Introduction
The existence of dark matter (DM) is a well-established paradigm in modern cosmology and astrophysics, supported by a wide range of observations, from galactic rotation curves to the anisotropies of the cosmic microwave background and the formation of large-scale structure. Yet, despite its dominant role in the matter content of the Universe, the particle nature of DM remains elusive Billard and others (2022).
In recent years, the field of direct DM detection has undergone significant advancements in sensitivity. However, no unambiguous signal has yet been observed. As the experimental landscape approaches the so-called neutrino fog O’Hare (2021), where Coherent Elastic Neutrino-Nucleus Scattering (CENS) Freedman (1974) becomes a dominant background, the need for diversified detection strategies becomes increasingly evident.
In this context, a multi-target and multi-detector approach is not only desirable but essential. Exploring different nuclear targets offers a powerful handle to disentangle potential DM signals from backgrounds, given the different dependencies of the interaction rates on nuclear parameters such as mass number and spin content Billard and others (2022). Furthermore, leveraging different detector technologies helps mitigate systematic uncertainties and enhances the robustness of any potential discovery.
Interestingly, the quest for direct DM detection shares deep experimental and conceptual similarities with the detection of low-energy neutrinos via CENS. Both processes involve low-energy nuclear recoils and require detectors with low energy thresholds, ultra-low background environments, and high radiopurity materials. As such, the development of detectors optimized for CENS can directly benefit DM searches, and vice versa Billard and others (2022).
In this paper, we report on a novel direct dark matter (DM) search that, for the first time, employs a PbWO4 cryogenic detector fabricated from archaeological Pb Pattavina and others (2019) and operated by the RES–NOVA collaboration. The RES–NOVA experiment is currently under commissioning and is primarily dedicated to the detection of low–energy neutrinos from astrophysical sources via coherent elastic neutrino–nucleus scattering Pattavina et al. (2020b). Thanks to its old age, the archaeological Pb used in this work is expected to exhibit an exceptionally low intrinsic radioactive background, making it particularly well suited for rare–event searches Pattavina et al. (2020a, b); Beeman and others (2013).
The results presented here represent an important step toward establishing PbWO4 as a competitive target material for both DM and coherent elastic neutrino–nucleus scattering (CENS) detection, exploiting its high mass number and favorable nuclear properties.
2 Archaeological Pb as a target material for elusive-rate searches
Direct searches for DM and the detection of low–energy neutrinos via CENS involve extremely low interaction rates, due to the small cross sections of these processes. Typical CENS cross sections lie in the range – Cadeddu et al. (2023), while for DM–nucleus elastic scattering current exclusion limits probe cross sections below Billard and others (2022) for DM masses in the GeV–TeV range. Consequently, the detection of such rare events requires detectors with exceptionally low background levels and very low energy thresholds.
To achieve this goal, particular care must be devoted to the selection and preparation of detector materials. One of the dominant background sources in rare-event searches is the intrinsic radioactivity of the materials used in the detector construction and its surrounding infrastructure Laubenstein and Lawson (2020). Long-lived isotopes from the uranium and thorium decay chains, as well as 210Pb and cosmogenic activation products, can generate signals in the energy region of interest and thus mimic or obscure the expected signal Clemenza et al. (2011).
In this context, archaeological Pb offers a unique advantage thanks to its extremely long cool-down time since production. Having been shielded from cosmic rays for centuries, typically underwater or underground, this material exhibits strongly suppressed levels of radioactive isotopes, in particular 210Pb and its progeny Pattavina and others (2019). Its intrinsic radiopurity makes archaeological Pb an excellent candidate for use in low–background detectors, either as passive shielding or as an active component, as in the case of PbWO4 crystals Pattavina and others (2021).
In the present work, we exploit the properties of archaeological Pb to develop and operate a cryogenic detector based on PbWO4 crystals. This approach allows us to investigate the potential of this material for direct DM detection, while simultaneously exploring its suitability for CENS searches. The combination of a high- target, a low energy threshold, and an ultra-low intrinsic background represents a promising pathway toward the observation of rare nuclear recoils associated with physics beyond the Standard Model Drukier and Stodolsky (1984).
A preliminary study of the sensitivity of PbWO4 to both spin-dependent and spin-independent DM interactions was presented in Ref. Alloni and others (2025). In that work, the RES-NOVA collaboration exploits the presence of both heavy (Pb) and light (O) nuclei to explore a wide range of DM masses spanning nearly four orders of magnitude. With an expected target mass of approximately 200 kg and a total exposure of 1 yr, RES-NOVA is projected to probe previously unexplored regions of the DM parameter space, providing complementary information to that obtained by established technologies such as Xe-based time-projection chambers Aprile and others (2023); Meng and others (2021); Aalbers and others (2023).
| Chain | Nuclide | Activity |
| [mBq/kg] | ||
| 232Th | 232Th | 0.04 |
| 228Th | 0.800.09 | |
| 238U | 238U | 0.03 |
| 234U | 0.03 | |
| 230Th | 0.04 | |
| 226Ra | 11.340.35 | |
| 210Pb/210Po | 22.500.49 | |
| 40K | 40K | 9 |
3 Experimental set-up
The crystal used in this work was obtained by cutting a 13 g sample from the same kg-scale PbWO4 crystal previously operated by the RES–NOVA collaboration in Ref. Beeman and others (2022). That large-mass crystal was specifically employed to assess the intrinsic radioactive background of the material, benefiting from the long accumulated exposure and the ultra–low–background conditions of the experimental setup.
In particular, the kg-scale crystal was operated within the cryogenic infrastructure that previously hosted the CUORICINO Arnaboldi and others (2008), CUORE–0 Alfonso and others (2015), and CUPID–0 Azzolini and others (2018) experiments, enabling a detailed characterization of its intrinsic radionuclide content. A summary of the measured radioactive contaminants is reported in Table 1.
The detector was operated in the Ieti dilution refrigerator Barucci et al. (2009), a custom-built cryostat designed for detector R&D in rare-event searches and optimized for mechanical stability and vibrational isolation. The Ieti cryostat provides a dry dilution refrigeration cycle, reaching a base temperature below 7 mK and offering a cooling power of approximately 500 W at 100 mK. Its compact design and reduced mechanical footprint make it ideal for both R&D and small-scale demonstrator experiments, despite the relative high-background operating conditions. This system is hosted in the Hall-C of the underground laboratory of Gran Sasso of INFN (Italy).
3.1 Cryostat Configuration
The cryostat has multiple thermal stages (50 K, 4 K, Still, Cold Plate, and Mixing Chamber) and a pulse tube refrigerator (PTR) for precooling, see Fig. 1. Each stage is thermally anchored to minimize temperature gradients and parasitic heat loads. The detector module is suspended from the Mixing Chamber (MC) plate using a low thermal conductivity structure engineered to reduce the transmission of mechanical vibrations, from the cryogenic infrastructure to the detector, while ensuring efficient thermal contact.
To further suppress the mechanical noise induced by the PTR, the cold head was mechanically decoupled from the cryogenic infrastructure and mounted on an independent support structure, physically isolated from the rest of the system. This configuration significantly reduces the transfer of vibrational noise to the detector volume.
The vibrational performance of the system was characterized using specially designed geophones operated at around 7 mK. These sensors enabled, for the first time, an in-situ assessment of vibrational noise within the actual operating environment of the detector (see Fig. 1). Previous characterizations Maisonobe et al. (2018); Wex and others (2025) reported in the literature rely exclusively on piezoelectric sensors operated at room temperature or near-to-room temperature, which do not provide a complete or reliable picture of vibrational noise propagation under actual experimental conditions. Indeed, the thermal contraction of a cryogenic system can reduce the overall dimensions of a system up to the cm–scale, impacting its intrinsic mechanical response de Wit et al. (2019). The use of cryogenic geophones in this set-up thus represents a significant advancement in understanding the impact of mechanical noise on cryogenic rare-event detectors.
In Fig. 2, the axial (vertical) vibrational level of the Ieti MC plate is shown in terms of the displacement amplitude spectral density (ASD), measured at a temperature of 7 mK. When compared with vibrational measurements performed in other cryogenic experimental setups Maisonobe et al. (2018); Wex and others (2025), the ASD indicates that the Ieti facility provides a particularly low–vibration environment, with vibrational amplitudes more than one order of magnitude lower in the low–frequency range (1–50 Hz), reaching the sub–nm level. This frequency range is particularly relevant for cryogenic detectors whose signal bandwidth covers this region.
3.2 Detector set-up
The detector consists of a PbWO4 crystal shaped as a rectangular crystal with dimensions of 0.7 0.7 4 cm3. The crystal is held in place and thermally coupled to the cryogenic environment by means of PTFE clamps, chosen for their low thermal conductivity and mechanical stability at mK temperatures. A Neutron Transmutation Doped (NTD) Ge thermistor Pirro and Mauskopf (2017) is glued to the crystal using Araldite epoxy resin, providing sensitive thermometric read-out of the heat pulses induced by particle interactions. The crystal is mounted inside an oxygen-free high-conductivity (OFHC) copper frame. In Fig. 3, a schematic view of the PbWO4 detector is shown.
The thermal signals from the NTD is read out via 25 m diameter gold bonding wires, which provide the necessary electrical connection while maintaining minimal thermal load. These wires deliver the signals to a contact pad array located on the MC stage. From this point, superconducting wires are used to transport the signals outside the cryogenic environment to the room-temperature read-out electronics, thereby minimizing parasitic heat leaks and preserving signal integrity.
The entire setup was operated underground at the Laboratori Nazionali del Gran Sasso (LNGS), benefiting from an overburden equivalent to 3600 m.w.e. Bellini and others (2012), which reduces the residual muon flux by approximately 6-orders of magnitude. The cryostat is housed within a passive shielding structure composed of non-archaeological Pb with a total thickness of 10 cm, surrounding the detector on the bottom and lateral sides. The top side of the detector is shielded by an additional 6 cm of non-archaeological Pb, arranged in two separate disks of 3 cm thickness each and installed directly inside the cryostat, one at the 4 K stage and one below the MC stage (see Fig. 1). The lateral sides are further surrounded by a 5 cm-thick Cu shield. This shielding configuration provides an effective suppression of environmental -ray backgrounds while maintaining mechanical compatibility with the cryogenic infrastructure. Nevertheless, residual background contributions are expected to originate from the shielding components themselves, which were not selected for radiopurity but rather optimized for shielding effectiveness and mechanical constraints.
3.3 Readout and Electronics
The thermistor was read out using a room-temperature setup. The bias was provided via a pair of high-value (1.3 G + 1.3 G) load resistors characterized by low noise. These large resistance values minimize their parallel noise contribution. The maximum bias voltage across the load resistors is V. This symmetric biasing circuit enables signal amplification through a differential voltage preamplifier—an improved version of the design in Ref. Arnaboldi et al. (2018). This preamplifier features very low series noise, specifically below at Hz and in the white noise region. Its parallel noise is approximately , well below that of the load resistors. When accounting for the thermal noise of the detector’s connecting links, the total white series noise at the preamplifier stage is approximately . The dedicated power supply system ensures high stability and negligible thermal drift for both front-end and detector biasing Carniti et al. (2016). The preamplifier is followed by a programmable gain amplifier (8-bit resolution) to optimize the signal range for the acquisition system. This system Carniti et al. (2023) consists of a frequency-programmable, 6-pole Bessel-Thomson antialiasing filter (ranging from Hz to kHz with 10-bit resolution) and a 24-bit - ADC. Data were acquired in continuous mode with a sampling frequency configurable up to kHz. The data stream was transmitted via a standard User Datagram Protocol (UDP) over an optical interface, enabling both online and offline processing.
For the measurements presented in this manuscript, the detector operating parameters were optimized to maximize the signal-to-noise ratio, following the same procedure described in Ref. Azzolini et al. (2018). The bias voltage was set to 22 V (corresponding to a bias current of 8.5 nA), the amplifier gain to 1030, and the sampling frequency to 10 kHz. At base temperature (7 mK) and under operating conditions (12 mK), the Ge–NTD thermistor exhibited resistance values of 36 M and 5.4 M, respectively.
4 Data Processing and detector performance
The continuous data stream, accounting for a total exposure of 32.4 gday, was processed through the newly developed analysis chain for RES–NOVA. The raw data stream was replicated into two parallel and synchronously processed branches. In the first branch, the signal was AC-coupled to remove the baseline offset and subsequently passed through a low-pass filter to suppress high-frequency noise. The filtered data were then processed in chunks of 2048 samples (204.8 ms), which were treated as candidate pulse windows, as shown in see Fig. 4. Each window was timestamped at the position of the maximum output of the filter, such that any particle-induced signal could be accurately located in the second, unfiltered data stream. Events extracted from this unfiltered branch were subsequently characterized using standard pulse-shape parameters.
This analysis pipeline operates in a trigger-less mode. When pulses are present, the filter output tags their position and amplitude; when no signal is present, the waveform is processed in the same way and timestamped at the filter’s maximum output within the 2048 samples. This approach allows the Region of Interest (RoI) to be defined a posteriori, without imposing any threshold or selection bias at the data-acquisition stage.
In the first stage of the analysis, the identified pulses were used to characterize the detector response in terms of signal shape and noise power spectral density. The signal template was obtained by averaging approximately 1000 particle-induced pulses of similar amplitudes around 500 keV. The noise power spectral density was computed in an analogous manner by averaging the squared Fourier components of baseline segments with the same duration as the pulse windows but containing no physical events. These two ingredients were then combined to construct the Optimum Filter Gatti and Manfredi (1986) (OF), which maximizes the signal-to-noise ratio under the assumption that the input waveform follows the expected signal shape and the stationary noise model.
In the second stage of the analysis pipeline, the OF replaced the generic low-pass filter, enabling optimized tagging of small pulses. In addition, pulses identified in the raw data stream—using the timestamps provided by the filtered branch—were fitted to a pulse model. The model consisted of an ideal signal template supplemented by a third-order polynomial to account for slow baseline drifts. Pulse amplitudes were extracted using a Maximum-Likelihood Estimation (MLE) procedure. In Fig. 4, we show two representative pulse traces at different energies: one close to the detector energy threshold (at 3 keV) and one at higher energy (at 8 keV). The pulse near the detector energy threshold exhibits a signal–to–noise ratio close to unity.
This pipeline yields two independent estimators for the amplitude of particle-induced pulses: the OF amplitude and the Maximum-Likelihood Estimator (MLE) amplitude. The OF amplitude typically provides the best resolution, as it explicitly maximizes the signal-to-noise ratio under the assumption of a known pulse shape and stationary noise. The MLE amplitude, on the other hand, is obtained by fitting each individual pulse to an idealized detector-response model and is therefore more sensitive to pathological or poorly reconstructed events. In particular, since MLE relies on a constrained minimization, deviations from the expected pulse shape—such as pile-up, glitches, or baseline instabilities—can lead to biased or unstable amplitude estimates.
A simple event–selection criterion was adopted: the amplitude obtained from the OF was required to be consistent, within a given tolerance, with the amplitude extracted from the pulse fit. This condition ensures that only well–reconstructed pulses, whose shapes are compatible with the expected detector response, are retained for the subsequent analysis. Several tolerance values were tested independently, namely 5%, 10%, 20%, and 30%. Since this parameter strongly influences the final sample of accepted events, its variation provides an estimate of the systematic uncertainty associated with the selection procedure. For the analysis presented here we adopted a tolerance of 10%.
The pulse amplitude ultimately used for the analysis is that provided by the OF, which offers the optimal energy resolution.
The measured pulse amplitudes were converted into deposited energy by identifying the 2615 keV line from 208Tl in the amplitude spectrum of a calibration measurement, which was used to anchor the absolute energy scale of the detector. As an independent validation of the calibration, we verified that the calibrated spectrum correctly reproduces the characteristic onset at 46 keV associated with the combined gamma and beta emissions from 210Pb, also in the background measurement (see Fig. 5). Thanks to these clear signatures in the energy spectrum we can compute the detector signal amplitude: 112 V/keV, a value that is in line with other high-performance cryogenic detectors Beeman and others (2013); Cardani et al. (2012); Barucci et al. (2019); Helis and others (2024).
In Fig. 5, the background energy spectrum of all accepted events, including interactions and nuclear recoils, is shown. The spectrum exhibits a steep rise at low energies, around , which originates from the acceptance of events with a signal–to–noise ratio close to unity. These events may include contributions of non-physical origin or signals whose pulse shape is inconsistent with the detector response model. Stricter selection criteria suppress the fraction of such noise–dominated events in the accepted sample, while looser criteria, conversely, lead to an increased acceptance of these events. At the same time, stricter criteria reduce the overall detection efficiency, whereas looser selections enhance it. This suppressing or enhancing effect is observable down to energies of about ; below this threshold, the impact of the selection criteria on the spectrum is no longer appreciable, indicating a signal–to–noise ratio below unity. The threshold for event identification is primarily determined by the signal sensitivity of the thermal sensors employed in this study, namely Ge–NTD thermistors. In comparison with other sensor technologies better suited for low–energy DM searches, such as transition edge sensors (TES) Pirro and Mauskopf (2017), significantly lower thresholds are expected.
The dominant contribution to the total counting rate in the DM RoI (2.5–10 keV) originates from background sources external to the crystal. These are primarily associated with the cryogenic infrastructure, such as Bremsstrahlung radiation from 210Bi present in the internal Pb shielding, and with environmental backgrounds, mainly due to ambient neutrons and -ray radiation, as previously discussed in Ref. Alloni and others (2025). Dedicated Monte Carlo studies will be performed to assess in more detail the different contributions. To further illustrate this point, Fig. 5 also shows the expected contribution from radioactive bulk contaminations within the operated PbWO4 crystal. This component is evaluated through Monte Carlo simulations of the experimental setup, using as input the results of the high-statistics radioactive screening measurements performed on the same crystal and summarized in Tab. 1.
To evaluate the detection and reconstruction efficiency in the RoI for DM investigations, we injected a comb of simulated mono-energetic pulses directly into the raw data stream and processed them through the full analysis pipeline. For each simulated energy, events that passed the quality selection and were correctly reconstructed were counted as successful detections. The efficiency , dependent on the chosen quality selection tolerance, was then obtained as the fraction of successfully reconstructed events over the total number of injected pulses at each energy value. In Fig. 6, the results of this analysis step are shown. In addition, for the evaluation of the detector baseline resolution, we fit a Gaussian distribution to the simulated lines obtaining: 234 eV, see Fig. 7.
5 Dark Matter search
In order to derive the sensitivity to DM particles, we compare the observed energy spectrum with the expected one in PbWO4 for a given DM particle mass and scattering cross-section. The expected recoil rate in the target crystal is computed using the wimprates package developed by the Xenon Collaboration Aalbers et al. (2023), including contributions from Pb, W, and O isotopes. For the spin-dependent interaction, we incorporate the nuclear spin structure of 207Pb, as discussed in Ref. Alloni and others (2025).
The DM halo model parameters are taken from Ref. Baxter and others (2021), assuming a local DM density of , an escape velocity of for the galaxy, a local standard-of-rest velocity of , a solar peculiar velocity of , and an average galactocentric Earth speed of . The recoil spectrum is subsequently folded with the detector response, accounting for the measured detection efficiency and energy resolution, in order to obtain the expected DM signal at the end of the analysis pipeline.
Although the experimental setup is installed underground, it is not optimized for ultra-low-background measurements. Moreover, Ge-NTD sensors do not provide the same performance as TESs for very small signal amplitudes, leading to a reduced efficiency of data selection at the lowest energies. As a consequence, no reliable background model can be assumed in the RoI. We therefore employ Yellin’s optimum interval method Yellin (2002) to derive a robust—albeit conservative—estimate of the sensitivity to the DM particle–nucleon scattering cross-section at 90% CL.
The lowest cross sections that can be excluded by this measurement, as a function of the DM particle mass, are shown in Fig. 8 for both spin–independent and spin–dependent interactions on 207Pb and 17O neutrons. In this evaluation, we neglect the contribution of 183W, which has the same nuclear spin as 207Pb, due to its lower natural isotopic abundance. These results represent the first constraints obtained using PbWO4 operated as a target material. The achieved sensitivity is currently limited by the detector exposure, amounting to only , and by the environmental background associated with the cryogenic infrastructure. Operation of the same detector in a lower–background environment would substantially extend the sensitivity of this approach in probing the DM parameter space.
6 Conclusions and perspectives
We have successfully operated a cryogenic detector prototype based on a PbWO4 crystal grown from archaeological Pb and established a robust data-analysis framework enabling the derivation of exclusion limits on DM interactions with ordinary matter, for the first time using this compound. Although the achieved sensitivity is not yet competitive with that of state-of-the-art experiments, this measurement provides a clear proof of concept of the detector technology. While not representative of the final RES–NOVA detector design, the prototype demonstrates effective control of mechanical noise in a cryogenic system at a level compatible with low-energy rare-event searches. Furthermore, the developed analysis methods —encompassing noise characterization, event reconstruction, and background handling — are readily transferable to Supernova neutrino detection via CENS. Together, these results represent a key validation step toward the scientific objectives of the RES–NOVA program.
The results presented in this work validate the use of PbWO4 crystals grown from archaeological materials in cryogenic rare-event searches. A significant improvement in sensitivity is expected from future implementations employing optimized low-background experimental setups and TESs in place of Ge-NTD thermistors, which are known to provide superior performance at low energies. In combination, these developments are expected to substantially lower the achievable energy threshold and enhance background discrimination capabilities.
In addition, we have validated the full data-analysis pipeline, which is capable of processing the detector data stream in real time. In the context of dark matter searches, this live processing enables continuous monitoring of detector performance and will be instrumental for operating-point stabilization and quality control once TES-based readout is implemented. In the context of Supernova neutrino detection, the same pipeline provides the capability to identify collective low-energy excesses and to issue alerts within seconds of a burst, thereby enabling prompt multi-messenger follow-up observations.
Acknowledgements.
The RES–NOVA collaboration thanks the directors and staff of the Laboratori Nazionali del Gran Sasso for their support, as well as the technical staff involved in the experiment. This work received funding from the European Union’s Horizon Europe programme through the ERC grant ERC–101087295 RES–NOVA. Additional financial support was provided under the National Recovery and Resilience Plan (NRRP) by the Italian Ministry of University and Research (MUR), funded by the European Union – NextGenerationEU, within the project “Advanced techniques for a next–generation isotopically enriched cryogenic Dark Matter experiment” (2022L2AXP2). This work made use of the Arby software for Geant4–based Monte Carlo simulations, developed within the framework of the Milano–Bicocca activities and currently maintained by O. Cremonesi and S. Pozzi. We gratefully acknowledge the University of Milano–Bicocca and INFN for their continuous support of the collaboration. S. Ghislandi acknowledges support from the U.S. Department of Energy (DOE) Grant No. [DE-SC0011091]. F.A. Danevich and V.I. Tretyak thank the Laboratori Nazionali del Gran Sasso, the Gran Sasso Science Institute, and the RES–NOVA Collaboration for their outstanding support and hospitality during the difficult period following the Russian invasion of Ukraine.References
- First Dark Matter Search Results from the LUX-ZEPLIN (LZ) Experiment. Phys. Rev. Lett. 131 (4), pp. 041002. External Links: 2207.03764, Document Cited by: §2, Figure 8, Figure 8.
- JelleAalbers/wimprates: v0.5.0 External Links: Document Cited by: §5.
- First results from the CRESST-III low-mass dark matter program. Phys. Rev. D 100 (10), pp. 102002. External Links: 1904.00498, Document Cited by: Figure 8, Figure 8.
- Low-Mass Dark Matter Search with the DarkSide-50 Experiment. Phys. Rev. Lett. 121 (8), pp. 081307. External Links: 1802.06994, Document Cited by: Figure 8, Figure 8.
- Search for Neutrinoless Double-Beta Decay of 130Te with CUORE-0. Phys. Rev. Lett. 115 (10), pp. 102502. External Links: Document, 1504.02454 Cited by: §3.
- New dark matter direct search based on archaeological Pb. Phys. Rev. D 111 (10), pp. 103050. External Links: 2501.12409, Document Cited by: §2, §4, §5.
- Light Dark Matter Search with Ionization Signals in XENON1T. Phys. Rev. Lett. 123 (25), pp. 251801. External Links: 1907.11485, Document Cited by: Figure 8, Figure 8.
- First Dark Matter Search with Nuclear Recoils from the XENONnT Experiment. Phys. Rev. Lett. 131 (4), pp. 041003. External Links: 2303.14729, Document Cited by: §2, Figure 8, Figure 8.
- A front-end electronic system for large arrays of bolometers. J. Instrum. 13 (02), pp. P02026. External Links: Document Cited by: §3.3.
- Results from a search for the 0 neutrino beta beta-decay of Te-130. Phys. Rev. C 78 (3), pp. 035502. External Links: 0802.3439, Document Cited by: §3.
- CUPID-0: the first array of enriched scintillating bolometers for 0$$nu beta beta $$decay investigations. Eur. Phys. J. C 78 (5), pp. 428. External Links: Document Cited by: §3.3.
- CUPID-0: the first array of enriched scintillating bolometers for decay investigations. Eur. Phys. J. C 78 (5), pp. 428. External Links: Document, 1802.06562 Cited by: §3.
- Cryogenic light detectors with enhanced performance for rare event physics. Nucl. Instrum. Methods Phys. Res. A 935, pp. 150–155. External Links: 1901.11009, Document Cited by: §4.
- A dry dilution refrigerator for the test of cuore components. J. Low Temp. Phys. 157 (5), pp. 541. External Links: Document Cited by: §3.
- Recommended conventions for reporting results from direct dark matter searches. Eur. Phys. J. C 81 (10), pp. 907. External Links: 2105.00599, Document Cited by: §5.
- New experimental limits on the alpha decays of lead isotopes. Eur. Phys. J. A 49, pp. 50. External Links: Document, 1212.2422 Cited by: §1, §4.
- Radiopurity of a kg-scale cryogenic detector produced from archaeological Pb for the RES-NOVA experiment. Eur. Phys. J. C 82 (8), pp. 692. External Links: Document, ISBN 1434-6052 Cited by: Table 1, Table 1, §3, Figure 5, Figure 5.
- Cosmic-muon flux and annual modulation in borexino at 3800 m water-equivalent depth. J. Cosmol. Astropart. Phys. (05), pp. 015. External Links: Document Cited by: §3.2.
- Direct detection of dark matter APPEC committee report. Rep. Prog. Phys. 85 (5), pp. 056201. External Links: Document Cited by: §1, §1, §1, §2.
- A view of coherent elastic neutrino-nucleus scattering. EPL 143 (3), pp. 34001. External Links: 2307.08842, Document Cited by: §2.
- Performance of a large TeO2 crystal as a cryogenic bolometer in searching for neutrinoless double beta decay. J. Instrum. 7, pp. P01020. External Links: 1106.0568, Document Cited by: §4.
- A low noise and high precision linear power supply with thermal foldback protection. Rev. Sci. Instrum. 87 (5), pp. 054706. External Links: ISSN 0034-6748, Document Cited by: §3.3.
- High resolution filtering and digitization system for cryogenic bolometric detectors. Nucl. Instrum. Methods Phys. Res. A 1045, pp. 167658. External Links: ISSN 0168-9002, Document Cited by: §3.3.
- Radon-induced surface contaminations in low background experiments. Eur. Phys. J. C 71 (11), pp. 1805. External Links: Document Cited by: §2.
- Vibration isolation with high thermal conductance for a cryogen-free dilution refrigerator. Rev. Sci. Instrum. 90 (1), pp. 015112. External Links: 1810.06847, Document Cited by: §3.1.
- Principles and Applications of a Neutral Current Detector for Neutrino Physics and Astronomy. Phys. Rev. D 30 (11), pp. 2295. External Links: Document Cited by: §2.
- Coherent Neutrino Nucleus Scattering as a Probe of the Weak Neutral Current. Phys. Rev. D 9, pp. 1389–1392. External Links: Document Cited by: §1.
- Processing the signals from solid-state detectors in elementary-particle physics. Riv. Nuovo Cim. 9 (1), pp. 1–146. External Links: Document Cited by: §4.
- First measurement of Gallium Arsenide as a low-temperature calorimeter. Eur. Phys. J. C 84 (7), pp. 749. External Links: 2404.15741, Document Cited by: §4.
- Low Background Radiation Detection Techniques and Mitigation of Radioactive Backgrounds. Front. in Phys. 8, pp. 577734. External Links: Document Cited by: §2.
- Vibration decoupling system for massive bolometers in dry cryostats. J. Instrum. 13 (08), pp. T08009. External Links: 1803.03463, Document Cited by: §3.1, §3.1.
- Dark matter search results from the PandaX-4T commissioning run. Phys. Rev. Lett. 127 (26), pp. 261802. External Links: Document Cited by: §2.
- New definition of the neutrino floor for direct dark matter searches. Phys. Rev. Lett. 127 (25), pp. 251802. External Links: Document Cited by: §1.
- Production and characterisation of a cryogenic detector from archaeological Pb. Eur. Phys. J. A 56 (2), pp. 38. External Links: Document Cited by: §1.
- Radiopurity of an archeological Roman Lead cryogenic detector. Eur. Phys. J. A 55, pp. 127. External Links: Document, 1904.04040 Cited by: §1, §2.
- RES-NOVA sensitivity to core-collapse and failed core-collapse supernova neutrinos. J. Cosmol. Astropart. Phys. 2021 (10), pp. 064. External Links: Document Cited by: §2.
- Neutrino observatory based on archaeological lead. Phys. Rev. D 102 (6), pp. 063001. External Links: 2004.06936, Document Cited by: §1.
- Advances in Bolometer Technology for Fundamental Physics. Ann. Rev. Nucl. Part. Sci. 67, pp. 161–181. External Links: Document Cited by: §3.2, §4.
- Decoupling pulse tube vibrations from a dry dilution refrigerator at milli-Kelvin temperatures. J. Instrum. 20 (05), pp. P05022. External Links: 2501.04471, Document Cited by: §3.1, §3.1.
- Finding an upper limit in the presence of unknown background. Phys. Rev. D 66 (3), pp. 032005. External Links: physics/0203002, Document Cited by: §5.